OpenStudy (anonymous):

heyy guyss.... how about this ... x^3 +6x^2-x+1 divided by ... x-1 .. division of polynomials ...... i expecting that you can help me ....

OpenStudy (mimi_x3):

(x-1)(x^2 + 7x+ 6)

OpenStudy (chaise):

Now you can cancel out your x-1's and you're left with x^2+7x+6 If you're solving for x then you make the equation equal to zero and you can use the quadratic formula. :)

OpenStudy (chaise):

You could also factorise your answer and make it equal to zero (this might be easier) you get: (x+6)(x+1) = 0 For this to be true (atleast one variable must equal 0) So: x+6 = 0 x = -6 x+1 = 0 x = -1

OpenStudy (anonymous):

thanks guysssss.... ......

OpenStudy (anonymous):

ammhhpff .. where did the number 6 came from ???

OpenStudy (anonymous):